Solve the full problem


X2uxx +2xyuxy+uyy=u


1
Expert's answer
2021-05-07T14:24:04-0400

The given equation is-

"x^2u_{xx} +2xyu_{xy}+u_{yy}=u"


Change of variable method-


From this eqn."b^2-4ac=4x^2y^2-4x^2y^2=0" so eqn is parabolic


General form-

"au_{xx}+2bu_{xy}+cu_{yy}+du=0"


Charcterstics equation is-

"\dfrac{dy}{dx}=\dfrac{b}{a}=\dfrac{2xy}{x^2}=\dfrac{2y}{x}"


Let "h=x ,t=\dfrac{y}{x}"


"\Rightarrow u_x=u_x(1)+u_t(\dfrac{-y}{x^2})=u_x-\dfrac{y}{x^2}u_t



u_y=u_h(0)+u_t(\dfrac{1}{x}=\dfrac{1}{x}u_t"


"u_{xx}=u_{hh}-\dfrac{2y}{x^2}u_{ht}+\dfrac{y^2}{x^4}u_{tt}+\dfrac{2y}{x^3}u_t~~~~~~~-(1)\\[9pt]u_{yy}=\dfrac{1}{x^2}u_{tt}



~~~~~~-(2)\\[9pt]u_{xy}=\dfrac{1}{x}u_{xt}-\dfrac{y}{x^3}u_{tt}-\dfrac{1}{x^2}u_t~~~~~~~~-(3)"



Substituting eqn. (1), (2) and (3) in given Pde and we get-


"h^2u_{hh}-2yu_{ht}+t^2u_{tt}+2tu_t+t^2u_{tt}+2yu_{ht}-2t^2u_{tt}-2tu_{t}=u"


"h^2u_{hh}=u"


"u_{hh}=\dfrac{u}{h^2}"


or "\dfrac{d^2u}{dh^2}=\dfrac{u}{h^2}"


Integrating and we get-


"logu\dfrac{du}{dh}=\dfrac{-1}{h}+f(t)"


Again Integrating and we get-


"u_{h,t}=-lnh+hf(t)+g(t)"


On substituting the values of h and t we get-


"u(x,y)=2x^2+xf(\dfrac{y}{x})+g(\dfrac{y}{x})"